Determine whether the equation has two distinct real number solutions, one real numbers solution, or no real number solution.
x^2 - 8x + 16 = 0

Found 2 solutions by checkley77, solver91311:
Answer by checkley77(12844)

x^2 - 8x + 16 = 0
(x-4)^2=0
x-4=0
x=4 ans.
One real number solution.

Answer by solver91311(24713)



Compute the discriminant. The discriminant is the part of the quadratic formula , that is under the radical, namely

Two real and unequal roots. If is a perfect square, then the roots are rational numbers, otherwise the roots are irrational.

One real root with a multiplicity of two. That is to say that the trinomial is a perfect square and has two identical factors.

No real solutions. Rather you have a conjugate pair of complex roots of the form where is the imaginary number defined by


John